What is the more likely sum of the numbers when throwing two dices: 9 or 10?
1
Expert's answer
2010-06-22T05:39:03-0400
Numbers 9 and 10 can be received in two different ways: 9 = 3 + 6 = 4 + 5 and 10 = 4 + 6 = 5 + 5. But we have to consider the order in which the numbers can appear when throwing the dices. Thus, number 9 can be received in four different ways and number 10 – in three: 9 = 3 + 6 = 6 + 3 = 4 + 5 = 5 + 4, and 10 = 4 + 6 = 6 + 4 = 5 + 5.